OpenStudy (mathmath333):

divya goes to buy fruits and after lot of bargaining is able to get the price of a dozen of apples reduced by rs. 1 from the initial price, thereby enabling her to get 1 apple extra for every rupee saved (getting no discount on extra apples ). what is initial price of a dozen of apple ?. a.11 b.12 c.13 d.15

OpenStudy (cwrw238):

Hint: after the price is reduced by 1 rupee the cost will be 1 rupee per apple

OpenStudy (mathmath333):

i know the answer , i want to form and solve through equations to get answer

OpenStudy (cwrw238):

If the original price of 12 apples is x rupees then the cost of 12 after the discount is (x - 1) / 12

OpenStudy (mathmath333):

u mean \[x=\frac{ x-1 }{ 12 }\]

OpenStudy (mathmath333):

note that she gets \(1\) apple extra for the discounted price and the extra apple is not discounted

OpenStudy (cwrw238):

yes I appreciate that now...

OpenStudy (cwrw238):

I meant that the cost of 1 apple after the discount is (x-1)/12

OpenStudy (cwrw238):

though I must admit I'm confused by this problem...

OpenStudy (mathmath333):

orignal price is \(x\) for 12 apples new price is \(x-1\) for 13 apples

ganeshie8 (ganeshie8):

12 is the answer from your original hint right ?

OpenStudy (cwrw238):

yes

OpenStudy (mathmath333):

yes 12 is right , but i m curious about the equations

ganeshie8 (ganeshie8):

1 rupee saved = 1 apple extra thats the only equation i see here :)

OpenStudy (mathmath333):

so u mean this cant be solved through equation approach

ganeshie8 (ganeshie8):

the question itself has the answer

ganeshie8 (ganeshie8):

`enabling her to get 1 apple extra for every rupee saved `

OpenStudy (mathmath333):

lol

ganeshie8 (ganeshie8):

:P

OpenStudy (cwrw238):

yes -its so obvious - somehow the more I thought about the more i got confused lol

OpenStudy (mathmath333):

i found a site where it said answer is \(13 \) http://testfunda.com/Answers/ViewQuestion.aspx?QID=9b5214f1-7577-490b-aa3d-a737f6537d2c

ganeshie8 (ganeshie8):

depends on how we interpret the problem but the problem clearly says the extra apple needs to be purchased without discount that means 1 extra apple = 1 rupee this yields 12 rupee per dozen for the actual price before discount

OpenStudy (cwrw238):

that's the crux of the matter - how you interpret the problem

OpenStudy (mathmath333):

lol english is more important in word problems

OpenStudy (cwrw238):

true

OpenStudy (asnaseer):

lets call the original price of a dozen apples \(p\). Then we know if we reduce this price by 1 rupee then each apple costs 1 rupee. Therefore:\[\frac{p-1}{12}=1\]Solve this equation to get original price of a dozen apples.

OpenStudy (mathmath333):

ur logic is also correct

OpenStudy (asnaseer):

From your statement above: \[x=\frac{x-1}{12}\]is incorrect

OpenStudy (mathmath333):

book says "\(\text{from the options u can work that if the orignal price }\\ \text{ was 12 rs per dozen ,the cost per apple would be rs 1.}\\ \text{if she is able to get a dozen of apples at a reduced price }\\ \text{(reduction of rs 1 per dozen), she would be able to purchase}\\ \text{1 apple extra for every rupee she saved}\)"

OpenStudy (asnaseer):

OK - I guess the key here is "divya goes to buy fruits and after lot of bargaining is able to get the price of a dozen of apples reduced by rs. 1 from the initial price" So she gets the price of a DOZEN apples reduced by rs 1 --> I INCORRECTLY then assumed the price of each apples would therefore be (p-1)/12

OpenStudy (asnaseer):

each apple STILL costs rs 1 It is just that she got a discount on buying a DOZEN

OpenStudy (asnaseer):

So @ganeshie8 has the correct logic to solve this question :)

OpenStudy (anonymous):

ghost riders again lol

OpenStudy (mathmath333):

i found the equation price= "p" real cost= 12 apples * "price p" = 12p discounted cost= 12 apples * (p-1) = 12(p-1) so, => real cost - discounted cost = an extra apple at price "p" => 12p - 12(p-1) =

OpenStudy (mathmath333):

\(12p-12(p-1)=p\)